please help.
solve the following systems of equations using substitution to eliminate one variable. if the system has no solution, indicate that it is inconsistent; if it has infinite solutions, indicate that it is dependant; if the system has one solution, check your answer.
y + 5x = 17
4x -3y = 6

Answer by mananth(16946)

5 x + 1 y = 17 .............1
-3 y = 6 -4 x ---------2
/ -3 =
y = -2 + 1.33 x
Substitute y in (1)
5 x + 1 ( -2 + 1.33 x ) = 17
5 x -2 + 1.33 x = 17
5 x + 1.33 x = 17 2
6.33 x = 19
/ 6.33
x= 3
Plug the value of x in (1)
5 x + 1 y = 17
5 * 3 + 1 y = 17
15 + 1 y = 17
1 y = 2
/ 1
y= 2
m.ananth@hotmail.ca